# Opinion de-polarization of social networks with GNNs

### v1

@@ 5 Algorithms
ML Based Approach
Now we discuss how the GCN can help derive a method to solve efficiently the problem ModerateExpressed. Intuitively, what we want to do is to train a GCN to predict the decrease of a polarization in the network, after setting a node’s external opinion to zero. This could, theoretically, be approximated from a GCN, because the embedding of a node can encode information about the structure and the opinions of its neighborhood. Thus, a GCN should be able to predict a high decrease in the polarization, if the embedding of the node $v$ that we consider to set $z_{v}=0$, suggests that node $v$ is a central node in a significantly polarized part of the network. Based on this intuition, we continue with our algorithm which is separated into two main components: the offline part and the online part.
@@
This part of our solution refers to the process of training the GCN to predict the decrease of the polarization after setting a node’s external opinion to zero. Let:
$Gain(v)=\pi(\mathbf{z_{old}})-\pi(\mathbf{z_{new}})$
where $\mathbf{z_{old}}$ and $\mathbf{z_{new}}$ are the vectors of external opinions before and after setting $z_{v}=0$ respectively.
In other words, we want to train a GCN to predict the $Gain$ for each node. This is a node regression task because it predicts a continuous target value. To be more precise, the GCN is responsible to generate the node embeddings. We then use those node embeddings as input to a regression head which outputs the predicted $Gain$. The regression head is usually a linear layer because the training is done in an end to end fashion and non-linearities are already applied in earlier stages of the pipeline (inside the GCN). In Section 6 we present the data that we used for the training of the GCN.
@@
The online part of our algorithm is what we actually use to solve the ModerateExpressed problem. It builds upon and aims to improve, in terms of time, the previously discussed GreedyExt algorithm. The core principle behind both algorithms is iterative: at each timestep, we aim to select the best node to add to the solution set based on its contribution to the reduction of polarization in the network. However, our approach introduces a key innovation that dramatically enhances efficiency and scalability.
@@
In summary, by integrating machine learning into the node selection process, our algorithm aims to significantly improve upon the GreedyExt method, in terms of time required to construct a solution set, that is almost as good as the one that GreedyExt finds. It maintains the iterative structure, but replaces brute-force computation with a more scalable approach, reducing complexity, while preserving accuracy in the identification of key nodes. We refer to this algorithm as the GNN Algorithm in the experiments that follow. Algorithm 2 contains the pseudocode for the GNN Algorithm.
Algorithm 2 GNN Algorithm
@@ 6 Datasets and Training
We used both synthetic and real networks, which we present in this section.
@@
Earlier we discussed about the node regression task, which we commonly refer to as the gain regression task. In order to train the GCN for this task we need data with ground truth targets. As a first step we construct synthetic graphs with two echo chambers. The graphs are generated using a degree corrected stochastic block model (DCSBM), which was introduced in (Karrer & Newman, 2011). The difference of this model with other stochastic block models lies in the degree distribution. In the DCSBM model the degree distribution of the graph follows a power law. The details of the DCSBM model can be found in the paper (Karrer & Newman, 2011).
Figure 2: Example of graph generated with the DCSBM model
@@
We construct 128 graphs with the DCSBM model, where each graph consists of 1000 nodes. The degrees are drawn from a power law distribution, which is common in social networks according to (Newman, 2010). Next we assign internal opinions to nodes based on their community membership. Nodes $v$ in community 1 are set $s_{v}=+1$ while nodes $u$ in community 2 are set $s_{u}=-1$. We don’t bother setting the external opinions, because their values at equilibrium don’t depend on their initial values. They depend on the vector $\mathbf{s}$ as Equation 2 suggests.
In order to use the generated graphs to train our model on the gain regression task, one more step is required. We have to compute the true $Gain$ for each node as the ground truth target value and then use it to train our model in a supervised setting. For this reason, the gain of each node has to be computed independently, as a preprocessing step for every graph of the dataset. Then the dataset can be used to train the model.
Figure 3: Retweet graph for the wiretaping scandal with 5k nodes
Our experiments suggested that it was beneficial to use simpler models in terms of the ModerateExpressed problem. The more complex models, although they performed better in terms of the gain regression task, they didn’t make as good node choices in the ModerateExpressed problem. This is probably because the more complex models could learn very well the properties of the graphs generated from the DCSBM model. In other words, we suspect that the complex models didn’t overfit the dataset but they overfitted the global properties of the synthetic graphs, which are naturally different than real graphs. However, at the end of the day we are interested in the ModerateExpressed problem and not in the gain regression. Consequently, larger loss in the gain regression task is acceptable, as long as the ranking of the nodes is correct. We are not interested in the exact value of the gain. We only want to be able to tell which node is the best choice at each timestep. Finally, a simpler model has an extra advantage in terms of the time required for the algorithm to complete. As a result, for our experiments we used a simple GCN with 2 layers where each layer consisted of 16 hidden units.
@@
We evaluated the GNN Algorithm in both real and synthetic networks. Here we present the real networks that we used. The experimental results are discussed in the next section.
@@
This dataset consists of political books that were sold on Amazon. The edges between political books represent frequent co-purchasing. The books are labeled as Conservative, Liberal and Neutral. The opinion mapping that we use is the following: Conservative: 1, Neutral: 0, Liberal: -1. The reason why we choose these 3 values is because we have no means to quantify numerically for each node what their exact opinion is. One alternative is to randomly sample opinions in the negative side of the spectrum of Liberal nodes, and in the positive side of the spectrum for the Conservative nodes. Finally, for the neutral nodes we could sample opinions around zero. There is no big difference in terms of our methodology. Of course, the resulting polarisation of the network and the final selected nodes by the algorithms would be different as their selection or not from the algorithms, depends both on their opinion and graph structure. But we are examining whether the GNN Algorithm that we propose makes equivalent choices to the GreedyExt algorithm, so the starting point of the network doesn’t really make any difference. In both cases the GreedyExt algorithm will make its greedy choices, and our algorithm should present similar behavior. For our experiments we decided to stick with the +1,0, -1 mapping to align with the equivalent experiments in (Matakos et al., 2017). The polarisation is $\pi=0.107$. It consists of 105 nodes and 441 edges.
@@ 7 Experiments
In this section we present the experimental results of our algorithm and compare it to the GreedyExt. For all experiments we set $K$ to be small, as allowing for large values makes the problem trivial. For example if we set $K$ to be equal to approximately 90% of the graph’s size, then we wouldn’t need to make sophisticated node choices, as most of the nodes would be by default set to zero and the polarization index would have been negligible. We limit $K$ to approximately 10% of the network’s size.
Additionally, in order to demonstrate that the problem is not trivial, we used a random algorithm as a baseline. The random algorithm picks $K$ nodes randomly and sets their external opinions to zero. Similar performance as the GreedyExt implies that the problem is trivial.
@@
Polarization
Figure 5 shows how the polarization changes on each step of the algorithms, for the Political Books dataset. For example the green point at $k=4$ denotes the polarization of the network after the GNN Algorithm has set to zero the first 4 nodes (out of $K$). We observe the following points. First of all, the random algorithm fails to decrease the polarization as much as the other algorithms, a fact which implies that the problem is not trivial. Additionally, the GNN Algorithm follows GreedyExt closely, which means that it makes equivalent choices at each timestep.
The results of our experiments for the real networks are summarized in Figure 6. According to this figure, the GNN Algorithm always achieves similar polarization to GreedyExt. The organe bar is the initial polarisation for each dataset, while the blue and the green bars show the polarization achieved from GreedyExt and GNN Algorithm respectively, after setting $K$ nodes to zero.
Figure 6: Polarization of each algorithm per dataset
Time
So far, we have only discussed how the two algorithms compare, in terms of the final polarization that they achieve. As we have already mentioned, the GreedyExt algorithm, is not efficient for large graphs, because of its high computational complexity. Particularly, the evaluation of every node independently at each timestep increases the complexity dramatically. In contrast, the GNN Algorithm replaces this brute-force style approach with a forward pass on a relatively small neural network, followed by an argmax operation. Figure 7 shows the required time for each algorithm to complete for the datasets LiveJournal and Wiretaping. For this experiment we set $K=50$.
@@
As the picture shows, the GNN Algorithm outperforms the GreedyExt in terms of time. Similar results hold for the other datasets as well. Finally, Figure 8 shows the required time for each algorithm to complete on graphs with various sizes. As we can see, the required time for the GreedyExt algorithm grows very quickly with respect to the graph size. In contrast the time to complete for the GNN Algorithm grows much slower, with respect to the size of the graph.
@@
Moreover, the time advantage of the GNN Algorithm against the GreedyExt, is larger, for larger graphs. In other words, it increases with respect to the size of the underlying graph. In the examples that we have shown where large graphs were involved, the LiveJournal and the Wiretaping datasets, the GNN Algorithm is ×$\times 10$ and ×$\times 16$ times faster respectively.
@@ 8 Conclusion
In this work, we proposed an alternative approach to address the ModerateExpressed problem, originally introduced in (Matakos et al., 2017). One of the key limitations of their solution is its lack of scalability, as their algorithm becomes increasingly slow, when applied to larger graphs. Our method leverages Graph Neural Networks (GNNs), specifically a 2-layer Graph Convolutional Network (GCN), to generate node embeddings. Instead of using brute force evaluation, we apply a simple argmax function to identify the node with the highest score. In terms of polarization, our approach achieves results almost identical to those of the original algorithm. However, the main advantage of our method is its ability to scale efficiently to larger graphs due to the GNN architecture. Our results show that the GNN Algorithm dramatically improves the required time to solve the problem and thus it is a great choice when dealing with large graphs. An interesting future direction is to explore the submodularity of the problem ModerateExpressed and quantify the suboptimality for greedy algorithms like the ones we have discussed.

### v2

@@ 5 Proposed Solution
GCN for ModerateExpressed
Now we discuss how the GCN can help derive a method to solve efficiently the problem ModerateExpressed. Intuitively, what we want to do is to train a GCN to predict the decrease of the polarization of a network, after setting a node’s external opinion to zero. This could, theoretically, be approximated from a GCN, because the embedding of a node can encode information about the structure and the opinions of its neighborhood. Thus, a GCN should be able to predict a high decrease in the polarization, if the embedding of the node $v$ that is being moderated, suggests that node $v$ is a central node in a significantly polarized part of the network. Based on this intuition, we continue with our method which is separated into two main components: the offline part and the online part.
@@
This part of our method refers to the process of training the GCN to predict the decrease of the polarization after setting a node’s external opinion to zero. Let:
$Gain(v)=\pi(\mathbf{z^{ss}_{old}})-\pi(\mathbf{z^{ss}_{new}})$
where $\mathbf{z^{ss}_{old}}$ and $\mathbf{z^{ss}_{new}}$ are the steady state vectors of external opinions before and after setting $z_{v}=0$ respectively.
In other words, we want to train a GCN to predict the $Gain$ for each node. This is a node regression task because it predicts a continuous target value. To be more precise, the GCN is responsible to generate the node embeddings. Then, those node embeddings are given as input to a regression head which outputs the predicted $Gain$. The regression head is usually a linear layer because the training is done in an end to end fashion and non-linearities are already applied in earlier stages of the pipeline (inside the GCN). We provide a detailed description of the data that were used for training in Section 6.
@@
The online part of our method is what we actually use to solve the ModerateExpressed problem. It builds upon and aims to improve, in terms of time, the previously discussed GreedyExt algorithm. The core principle behind both algorithms is iterative: at each timestep, we aim to select the best node to add to the solution set based on its contribution to the reduction of polarization in the network. However, our approach introduces a key innovation that dramatically enhances efficiency and scalability.
@@
In summary, by integrating machine learning into the node selection process, our algorithm aims to significantly improve upon the GreedyExt method, in terms of time required to construct a solution set, that is almost as good as the one that GreedyExt finds. It maintains the iterative structure, but replaces brute-force computation with a more scalable approach, reducing complexity, while preserving accuracy in the identification of key nodes. We refer to this algorithm as the GNN-GreedyExt in the experiments that follow. Algorithm 2 contains the pseudocode for the GNN-GreedyExt.
Algorithm 2 GNN-GreedyExt
@@ 6 Datasets and Training
We used both synthetic and real networks, which we present in this section. Realistic synthetic networks are required in order to facilitate the training process of the GCN [^4] It was in general not easy to find hydrated instances of social networks with two echo chambers and ground truth communities (red blue). Due to this limited availability of data we decided to train on artificial networks which try to capture some important properties of real networks, while real networks are necessary to evaluate our method on real scenarios.
@@
Earlier we discussed about the node regression task, which we commonly refer to as the gain regression task. In order to train the GCN for this task we need graphs where each node is accompanied with a target value that our model has to learn to predict. This value is no other than the gain that results from moderating the node (setting it’s expressed opinion to zero). As a first step towards this direction, we construct synthetic graphs with two echo chambers. The graphs are generated using a degree corrected stochastic block model (DCSBM), which was introduced in (Karrer & Newman, 2011). The difference of this model with other stochastic block models lies in the degree distribution. In the DCSBM model the degree distribution of the graph follows a power law. The details of the DCSBM model can be found in the paper (Karrer & Newman, 2011).
Figure 2: Example of graph generated with the DCSBM model.Node size reflects degree, with larger nodes having higher degrees. Blue and red nodes represent two echo chambers.
@@
We construct 128 graphs with the DCSBM model, where each graph consists of 1000 nodes. The degrees are drawn from a power law distribution, which is common in social networks according to (Newman, 2010). Next we assign internal opinions to nodes based on their community membership. Nodes $v$ in community 1 are set $s_{v}=+1$ while nodes $u$ in community 2 are set $s_{u}=-1$. Finally, $\mathbf{z^{ss}}$ is computed and $z_{i}^{ss}$ is attached to each node $i$. As a result every node consists of two features: $s_{i}$, and $z_{i}$.
As we mentioned earlier, training for the node regression task requires the actual gain resulting from the moderation of each node, as supervision. This is computed as a preprocessing step of our graph data.
Regarding the precise GCN architecture, we used a 2-layer GCN with 16 hidden units per layer resulting in 16 dimensional node embeddings. A linear layer with 16 input values and one output value follows to transform the node embeddings into predicted gains.
This is a relatively simple model with a small number of parameters, but it works very well in practice. Although more complex models with a large number of parameters perform slightly better in the node regression task, when it comes to inference (solve ModerateExpressed problem) we observe no difference in the solutions. This is probably because the inference (online) task depends on the ordering of the predicted nodes’ gains rather than their exact values. Simply put, despite the predicted gains of the simple model (16 dim - 2 layer GCN) being less accurate compared to complex models, the ordering of the simple model’s predictions is consistent with the ordering of the predictions of complex models (i.e. if node A is better than B based on the complex models’ predictions, the same holds for the simple model). For this reason the inferred nodes are the same in both cases.
Finally, a simpler model has an extra advantage in terms of the time required for the algorithm to complete compared to scenarios with more complex models. This is because a smaller neural network results in faster forward passes. As a result, for our experiments we used this simple GCN with 2 layers where each layer consisted of 16 hidden units.
@@
We evaluated the GNN-GreedyExt in both real and synthetic networks. Here we present the real networks that we used. The experimental results are discussed in the next section. Some important properties characterizing both synthetic and real networks are summarized in Table 1
Graph $|V|$ $|E|$ $\pi$
Books 105 441 0.107
LiveJournal 2766 24138 0.179
WireTaping 5000 201876 0.011
DCSBM 977 3724 0.102
Table 1: Characteristics of the real networks considered. The last row refers to a synthetic (DCSBM-generated) graph that was used for training
@@
This dataset consists of political books that were sold on Amazon. The edges between political books represent frequent co-purchasing. The books are labeled as Conservative, Liberal and Neutral. The opinion mapping that we use is the following: Conservative: 1, Neutral: 0, Liberal: -1. The reason why we choose these 3 values is because we have no means to quantify numerically for each node what their exact opinion is. One alternative is to randomly sample opinions in the negative side of the spectrum of Liberal nodes, and in the positive side of the spectrum for the Conservative nodes. Finally, for the neutral nodes we could sample opinions around zero. There is no big difference in terms of our methodology. Of course, the resulting polarisation of the network and the final selected nodes by the algorithms would be different as their selection or not from the algorithms, depends both on their opinion and graph structure. But we are examining whether the GNN-GreedyExt that we propose makes equivalent choices to the GreedyExt algorithm, so the starting point of the network doesn’t really make any difference. In both cases the GreedyExt algorithm will make its greedy choices, and our algorithm should present similar behavior. Thus, without loss of generality, we stick with the +1,0, -1 mapping that we mentioned which is consistent with (Matakos et al., 2017) [^5] This opinion mapping was used for this dataset in their work and facilitates comparison with their results. The polarisation for the Political Books dataset is $\pi=0.107$. It consists of 105 nodes and 441 edges.
@@
Figure 4: Retweet graph for the wiretaping scandal with 5k nodes
@@ 7 Experiments
In this section we present the experimental results of our algorithm and compare it to the GreedyExt. For all experiments we choose small values of $K$ (10% of the graph’s size)
In order to ensure that the problem instance we are dealing with is not “trivial”, we also implement a random baseline algorithm, that picks the $K$ nodes to moderate uniformly. Clearly, scenarios in which the random scheme is competitive to the GreedyExt one suggest that either (i) this scenario has not enough social “structure” with all nodes being more or less equivalent (in their influence role) and/or (ii) $K$ is too large for this scenario [^6] It is easy to see that as K increases, the problem of “correctly” picking nodes becomes easier.
@@
De-polariszation achieved
In this first set of experiments, we use one of our real datasets (the Politcal Books one) to demonstrate these two key points: (i) the discussed impact of $K$, but that the problem is not “trivial” for a range of such values, even in a small scenario, and (ii) that our scheme’s performance is close to the “optimal” GreedyExt one.
In Figure 5 we show the final polarization index of Equation 3, on the y axis, as the number of nodes $K$ to moderate increases, for three algorithms: (i) GreedyExt, (ii) GNN-GreedyExt and (iii) the random baseline.
Key Observations
  1. 1.
    The GNN-GreedyExt algorithm follows GreedyExt closely, which means that it makes equivalent choices at each timestep.
  2. 2.
    The random baseline algorithm fails to decrease the polarization as much as the other algorithms, a fact which implies that the problem is not trivial
Similarly, we evaluate our method on the other real networks. The results of our experiments for the real networks are summarized in Figure 6. According to this figure, the GNN-GreedyExt always achieves similar polarization to GreedyExt. The orange bar is the initial polarization for each dataset, while the blue and the green bars show the polarization achieved from GreedyExt and GNN-GreedyExt respectively, after setting $K$ nodes to zero. Precisely, the polarization decrease of GNN-GreedyExt is roughly 95% of that achieved by GreedyExt.
Figure 6: Polarization of each algorithm per dataset. The orange bars show the initial polarization of each network. The blue and green bars show the final polarization that GreedyExt and GNN-GreedyExt achieve
Inference speed
So far, we have only discussed how the two algorithms compare, in terms of the final polarization that they achieve. As we have already mentioned, the GreedyExt algorithm, is not efficient for large graphs, because of its high computational complexity. Particularly, the evaluation of every node independently at each timestep increases the complexity dramatically. This is exactly where we expect the GNN-GreedyExt to excel, achieving the demonstrated close to optimal depolarization much faster. In the following set of experiments we demonstrate two key points: (i) the GNN-GreedyExt outperforms the GreedyExt in terms of inference speed, and (ii) the inference speed advantage of GNN-GreedyExt increases as the depolarizing graph’s size increases.
In Figure 7 the two algorithms are compared over the LiveJournal and Wiretaping datasets. Each bar corresponds to the time required (in seconds) for the corresponding algorithm to complete. For this experiment we set $K=50$ without loss of generality.
@@
As the picture shows, the GNN-GreedyExt outperforms the GreedyExt in terms of time. Similar results hold for the other datasets as well.
Finally, the next experiment compares the time to complete of the two algorithms for graphs with various sizes. In Figure 8 the y-axis shows the required time to complete, while the x-axis has the graphs’ sizes. As we can see, the required time for the GreedyExt algorithm grows very quickly with respect to the graph size. In contrast the time to complete for the GNN-GreedyExt grows much slower, with respect to the size of the graph.
@@
Moreover, the time advantage of the GNN-GreedyExt against the GreedyExt, is larger, for larger graphs. In other words, it increases with respect to the size of the underlying graph. In the examples that we have shown where large graphs were involved, the LiveJournal and the Wiretaping datasets, the GNN-GreedyExt is ×$\times 10$ and ×$\times 16$ times faster respectively.
@@ 8 Conclusion
In this work, we proposed an alternative approach to address the ModerateExpressed problem, originally introduced in (Matakos et al., 2017). One of the key limitations of their solution is its lack of scalability, as their algorithm becomes increasingly slow, when applied to larger graphs. Our method leverages Graph Neural Networks (GNNs), specifically a 2-layer Graph Convolutional Network (GCN), to generate node embeddings. Instead of using brute force evaluation, we apply a simple argmax function to identify the node with the highest score. In terms of polarization, our approach achieves results almost identical to those of the original algorithm. However, the main advantage of our method is its ability to scale efficiently to larger graphs due to the GNN architecture. Our results show that the GNN-GreedyExt dramatically improves the required time to solve the problem and thus it is a great choice when dealing with large graphs. An interesting future direction is to explore the submodularity of the problem ModerateExpressed and quantify the suboptimality for greedy algorithms like the ones we have discussed.
Acknowledgements
The authors have been supported in part by the Hellenic Foundation for Research & Innovation (HFRI) project 8017: “AI4RecNets: Artificial Intelligence (AI) Driven Co-design of Recommendation and Networking Algorithms”